SPECTROSCOPY AND DYNAMICS OF I_{2}-Ne^{a}

\underline{GALINA KERENSKAYA}, AMY BURROUGHS, MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

The spectroscopy and dynamics of I_{2}-Ne have been re-examined using OODR and fluorescence depletion techniques. Action spectra for I_{2}(B, ν)-Ne, detected by monitoring the I_{2}(B, ν-1) predissociation fragments, show that the Δν=-1 predissociation channel is suppressed for levels above ν=33, and closed for ν>36. From these data we obtained a revised estimate for the dissociation energy for I_{2}(B)-Ne of D_{0}=57.5 cm^{−1}. Action spectra for I_{2}(B, ν=34)-Ne, detected by monitoring I_{2}(B, ν=33) fragments, revealed a progression of nine intermolecular vibrational levels that had not been observed previously. These levels have been assigned to T-shaped or line ar geometries of I_{2}(B, ν=34)-Ne. Assignments were based on characteristic vibrational distributions exhibited by the I_{2}(B, ν-Δν) predissociation fragments. Fluorescence depletion measurements show that all of the bands in the action spectrum originate from a common ground state level. Furthermore, the one atom cage effect fluorescence from I_{2}(B)-Ne can be depleted by transitions from the zero-point level of I_{2}(X)-Ne. These observations show that the ground state wa vefunction is delocalized, sampling both the T-shaped and linear configurations of the complex.

aWork supported by National Science Foundation.